# Pilot-Quantum: A Quantum-HPC Middleware for Resource, Workload and Task Management

### v2

@@ Abstract
As quantum hardware continues to scale, managing the heterogeneity of resources and applications – spanning diverse quantum and classical hardware and software frameworks – becomes increasingly critical. Pilot-Quantum addresses these challenges as a middleware designed to provide unified application-level management of resources and workloads across hybrid quantum-classical environments. It is built on a rigorous analysis of existing quantum middleware systems and application execution patterns. It implements the Pilot Abstraction conceptual model, originally developed for HPC, to manage resources, workloads, and tasks. It is designed for quantum applications that rely on task parallelism, including: (i) Hybrid algorithms, such as variational approaches, and (ii) Circuit cutting systems, used to partition and execute large quantum circuits. Pilot-Quantum facilitates seamless integration of quantum processing units (QPUs), classical CPUs, and GPUs, while supporting high-level programming frameworks like Qiskit and Pennylane. This enables users to design and execute hybrid workflows across diverse computing resources efficiently. The capabilities of Pilot-Quantum are demonstrated through mini-applications – simplified yet representative kernels focusing on critical performance bottlenecks. We present several mini-apps, including circuit execution across hardware and simulator platforms (e. g., IBM’s Eagle QPU), distributed state vector simulation, circuit cutting, and quantum machine learning workflows, demonstrating significant scale (e. g., a 41-qubit simulation on 256 GPUs) and speedups (e. g., 15x for QML, 3.5x for circuit cutting).
@@ I Introduction
Motivation: Quantum computing promises advancements in computational challenges across science and industry, with applications in chemistry, simulations, optimization, and machine learning [1]. Quantum algorithms leverage quantum information principles to outperform classical methods with fewer computational steps [2, 3]. While Noisy Intermediate-Scale Quantum (NISQ) hardware [4] has progressed, further improvements in qubit fidelity [5], error correction [6], and software are needed for a practical advantage [7].
@@
Quantum-HPC Integration: Quantum computing complements HPC by addressing computational challenges currently intractable. HPC supports quantum research by enabling scalable classical simulations essential for many hybrid quantum-classical algorithms. Tighter integration of classical and quantum resources will become increasingly important [8, 9, 10, 11, 12]. This integration presents challenges across several levels: (i) Hardware Integration: Efficiently combining heterogeneous quantum systems (e.g., superconducting and ion-trap qubits) with classical computing resources (e. g., GPUs and CPUs) to create cohesive hybrid systems. (ii) Software Ecosystems: Developing robust software frameworks, tools, and middleware systems that seamlessly integrate with existing HPC tools (e. g., CUDA and MPI runtimes). Achieving this integration is crucial for advancing quantum computing and unlocking its full potential in solving complex computational problems.
Challenges: In particular, software tools [13, 14, 15] and middleware [16, 17, 18] will become critical for hardware development. The fragmented quantum software stack, still in its infancy, requires integration across multiple layers – from qubit control and circuit execution to application workflows. NISQ and FTQC algorithms rely heavily on quantum-classical integration [19, 20]. With maturing algorithms, end-to-end application workflows integrating data acquisition, pre-processing, and post-processing will gain importance [21]. Resource, workload, and task management are critical as these workflows use heterogeneous resources. Coordinating quantum and classical resources, managing tasks across diverse infrastructures, and integrating multiple programming paradigms complicate quantum application scaling.
Quantum-HPC Middleware: We propose Pilot-Quantum, a middleware for resource, workload, and task management in quantum-HPC systems. It is based on the Pilot-Abstraction [22, 23] and a conceptual quantum-HPC middleware architecture [20], designed from an analysis of quantum-HPC execution patterns (motifs [19]). The pilot abstraction uses placeholder jobs to reserve and manage resources across heterogeneous infrastructures (e. g., HPC and clouds). It supports loosely coupled parallelism in quantum applications, e. g., variational algorithms and circuit cutting.
Pilot-Quantum enables resource management and scheduling at the application level. It integrates seamlessly with quantum software ecosystems, supporting both high-level frameworks like Qiskit and Pennylane and low-level frameworks like cuQuantum for hybrid workloads [19]. With application-level scheduling, applications can optimize resource allocation, essential for costly QPUs, enabling efficient resource use for diverse classical and quantum task-based workloads.
@@
This paper is structured as follows: We explore quantum software and middleware state-of-the-art in section II. The primary contribution is Pilot-Quantum, an open-source middleware addressing resource, workload, and task management challenges is introduced in section III. Section IV demonstrates its capabilities by extending quantum mini apps [19] to new motifs, including distributed state vector simulation, circuit cutting, and quantum machine learning (QML). We conclude with a qualitative comparison of quantum middleware systems with Pilot-Quantum and future work in section V.
@@ II Background and Related Work
This section analyzes quantum-classical software and middleware using a conceptual architecture model [20] and quantum-HPC integration modes and execution motifs [19].
@@ II-A Conceptual Middleware Architecture
The conceptual quantum-HPC middleware architecture [20] consists of four distinct layers, denoted as $L1$, $L2$, $L3$, $L4$: $L4$ is the workflow layer, which encapsulates application semantics and logical dependencies between the workflow stages and offers a high-level abstraction for developing the quantum-HPC workflows. The workload layer $L3$ manages the workload, a set of interdependent, quantum, and classical tasks emitted by $L4$. It optimizes the workload execution by partitioning tasks and assigning them to the available resources utilizing parallelism (e.g., ensemble and task parallelism).
@@ II-B Quantum-HPC Integration Modes and Motifs
Managing quantum-HPC workloads and determining quantum-classical task dependencies requires identifying integration modes and motifs (i. e., recurring execution patterns) between quantum-HPC components [19].
@@
are categorized as Quantum-x-HPC or HPC-x-Quantum based on the strength of coupling between quantum and classical tasks [8, 20]. Three modes are identified: HPC-for-Quantum, with strong coupling optimizing quantum-classical interactions (e. g., error mitigation [24], error correction [25, 26], dynamic circuits [27]); Quantum-in-HPC emphasizes loosely-coupled parallelism between classical and quantum resource found in variational quantum algorithms (VQA) [28] to accelerate HPC application; and Quantum-about-HPC, loosely coupling quantum elements into HPC applications for pre- and post-processing [29, 30].
@@ II-C Quantum Software Layers
frameworks typically provide quantum circuit model-based abstraction for describing quantum programs. Examples are Qiskit [14, 31], PennyLane [15], Cirq [32], low-level frameworks like CUDA-Q and AWS Braket [33], and high-level programming libraries, such as, Qrisp [34], Silq [35], and Q# [36]. At the $L4+$ levels, domain-specific programming libraries include Qiskit Optimization [37], Qiskit Nature [38], and OpenFermion [39].
@@
like Qiskit Runtime [40] and Braket Hybrid Jobs [41, 42] provide session and decorator abstractions, respectively, to co-allocate heterogeneous resources for hybrid quantum-classical applications. However, new execution patterns, such as the combination of quantum and machine learning capabilities, require new abstractions, including auto differentiation of quantum circuits (e. g., using PennyLane [15]) or training classical models using quantum task outputs [43]. Some frameworks like XACC [44], QCOR [45], and CUDA-Q [18] use a kernel-based programming model with compile toolchains to manage quantum-classical function interactions. For classical simulation of quantum systems, acceleration frameworks, e. g., cuQuantum [46] are critical for performance and scale.
@@
At the $L1$ layer, many quantum frameworks provide device abstractions and plugins for accessing quantum devices, such as Qiskit backends [47], Qiskit Runtime [40], and PennyLane plugins [48]. Intermediate representations like OpenQASM [49] and QIR [50, 51] aim to bridge the gap between hardware platforms and high-level programming constructs. Transpiler and compiler are used to transform a circuit to a hardware-optimized form and are provided by full-stack frameworks like Qiskit and specialized frameworks like Tket [52], BQSKit [53], and Staq [54] transpilers.
@@
Resource management and scheduling in quantum-HPC environments occur across multiple levels and time scales [9, 55], involving integration of coarse-grained job-level allocation (e. g., GPUs) with QPU allocations and fine-grained, application-level scheduling [23, 56]. Existing HPC resource management frameworks often assume exclusive resource access, leading to resource fragmentation and underutilization [57]. As algorithms scale across heterogeneous resources (e. g., CPUs, GPUs, QPUs), such resource management will result in suboptimal performance and utilization. Pilot-Quantum specifically addresses the need for an application-level resource management balancing system and application objectives.
@@
at the task and workload layer ($L3$ and $L4$) are needed to orchestrate quantum workloads across integrated quantum-HPC systems. QML workloads, for instance, involve heterogeneous classical-quantum tasks for gradient computation, backpropagation, optimization, and circuit execution, hence requiring a systematic approach to workload and task management.
@@ II-D Middleware for Resource, Workload and Task Management
Many quantum middleware systems emerged recently including, QFaaS [58], Tierkreis [59], Qibo [60], Quantum Framework (QFw) [61], Quantum Brilliance [62], Rasqal [63] and Zapata’s Orquestra [64]. We particularly investigate Qiskit Serverless [16], Covalent [17], and CUDA-Q [18] systems system, which have different design objectives and principles.
II-D1 Qiskit Serverless
[65, 66, 16, 67] serves as middleware for Qiskit, managing resources and workloads in hybrid quantum-classical infrastructures while supporting long-running applications and scaling across distributed resources (CPU, GPU, QPU). It extends cloud-based quantum hardware access to classical computing environments at the $L2$ and $L3$ layers, utilizing Apache Ray [68] to execute hybrid workloads.
@@ II-D1 Qiskit Serverless
II-D2 Covalent
[17, 71] orchestrates distributed computing workflows across Cloud-HPC resources. Its SDK provides $L4$ level abstraction using electron and lattice decorators for defining tasks and workflows (i. e., DAG of electrons), where quantum tasks are defined using qelectron decorator. The Dispatch service at the $L3$ layer accepts workloads (i. e., ready-to-run tasks), maps them to suitable resources, and dispatches them to appropriate backends.
@@ II-D3 CUDA-Q
[18] supports tight integrations within heterogeneous computing environments. It employs an LLVM compiler toolchain [74] to convert frontend code in C++ or Python to different heterogeneous backends, such as CPUs, CUDA-enabled GPUs, and QPUs. It utilizes distinct intermediate representations for each processor type (e. g., QIR for QPUs and NVVM IR for GPUs), allowing for targeted optimizations. The CUDA-Q runtime, utilizes MPI for parallel processing and CUDA as a low-level runtime. While particularly suitable for HPC-for-Quantum tight integration, CUDA-Q also supports Quantum-for-HPC modes, e. g., for VQAs. At $L4$, it provides a circuit-based programming model and a kernel API for offloading circuits to hardware or simulated QPUs. Application-level task management uses an MPI and CUDA-Runtime on the $L2$/$L1$ level.
@@ III Pilot-Quantum Design and Architecture
Pilot-Quantum enhances resource, workload, and task management across heterogeneous classical and quantum infrastructures for quantum applications and workflows. It is an implementation of the Pilot Abstraction [22, 23], a conceptual abstraction for application-level scheduling [22, 75]. It uses a placeholder job concept for application-level control of allocated resources, crucial for scaling heterogeneous workloads.
Abstraction for Resource Management: Pilot-Quantum’s abstractions enable applications to effectively decompose workflows, simplify scaling, and manage diverse task-level parallelism. It can efficiently parallelize and execute operations like circuit ensembles or data-parallel QML training asynchronously and concurrently.
Easy Deployment: Pilot-Quantum simplifies workflow deployment across distributed, heterogeneous infrastructures, including non-local quantum cloud environments, HPC MPI/CUDA environments for distributed state vector simulation, and loosely coupled workflow systems. It can also integrate with resource-specific optimization tools, such as QPU-specific transpilers.
@@ III-A Requirements and Design Objective
Infrastructure Heterogeneity: The system must address both software and hardware heterogeneity, encompassing a range of QPUs such as superconducting qubits, ion traps, neutral atoms, and simulated QPUs utilizing MPI or CUDA. This requires compatibility with diverse frameworks and tools, including transpilers, compilers, and runtimes. Examples include various simulator types (e. g., accelerated state vector and tensor network simulations leveraging MPI or CUDA) and machine learning runtimes like PyTorch and JAX.
@@
Resource management in quantum-classical computing is challenging due to diverse resource types and application demands. Dynamic resource allocation during workflow execution is crucial to address application and infrastructure dynamism when handling varied tasks and dependencies across heterogeneous platforms. This requires employing multi-level scheduling coordination with efficient resource allocation, workload partitioning, and task management strategies to further optimize application performance and system utilization.
Abstractions: Integrating quantum computing into HPC workflows requires user-friendly, flexible abstractions that hide quantum complexity while supporting various parallelism forms (e.g., task parallelism for circuits, distributed state vector simulations using MPI and cuQuantum, and differentiable computing for QML). These should facilitate efficient pipeline creation for end-to-end quantum applications with pre- and post-processing integrations. The abstractions must offer seamless interoperability with quantum and HPC infrastructures, further ensuring compatibility with frameworks like CUDA for GPU-accelerated tasks and MPI for distributed computing.

### v3

# Pilot-Quantum: A Middleware for Quantum-HPC Resource, Workload and Task Management

@@ Abstract
As quantum hardware advances, integrating quantum processing units (QPUs) into HPC environments and managing diverse infrastructure and software stacks becomes increasingly essential. Pilot-Quantum addresses these challenges as a middleware designed to provide unified application-level management of resources and workloads across hybrid quantum-classical environments. It is built on a rigorous analysis of existing quantum middleware systems and application execution patterns. It implements the Pilot Abstraction conceptual model, originally developed for HPC, to manage resources, workloads, and tasks. It is designed for quantum applications that rely on task parallelism, including (i) Hybrid algorithms, such as variational approaches, and (ii) Circuit cutting systems, used to partition and execute large quantum circuits. Pilot-Quantum facilitates seamless integration of QPUs, classical CPUs, and GPUs, while supporting high-level programming frameworks like Qiskit and Pennylane. This enables users to efficiently design and execute hybrid workflows across diverse computing resources. The capabilities of Pilot-Quantum are demonstrated through mini-apps – simplified yet representative kernels focusing on critical performance bottlenecks. We demonstrate the capabilities of Pilot-Quantum through multiple mini-apps, including different circuit execution (e.g., using IBM’s Eagle QPU and simulators), circuit-cutting, and quantum machine learning scenarios.
@@ I Introduction
Motivation: Quantum computing promises to advance specific science and industry computational challenges, from chemistry, optimization and simulations to machine learning [1], by theoretically outperforming classical methods by requiring fewer computational steps [2, 3]. While quantum hardware has progressed, e. g., concerning qubit fidelity [4] and error correction [5], further improvements are needed for a practical quantum advantage [6].
@@
Quantum-HPC Integration: To realize the benefits of quantum computing, tighter integration of classical and quantum resources will become increasingly important [7, 8, 9, 10, 11]. Such integration presents challenges across several levels: (i) Hardware Integration: Efficiently combining heterogeneous quantum systems (e.g., superconducting and ion-trap qubits) with classical computing resources (e. g., GPUs and CPUs) to create cohesive hybrid systems. (ii) Software Ecosystems: Developing robust software frameworks, tools, and middleware systems that seamlessly integrate with existing HPC tools (e. g., CUDA and MPI runtimes). Achieving this integration is crucial for advancing quantum computing and unlocking its full potential in solving complex computational problems.
Challenges: Resource, workload, and task management in classical workflow systems face significant challenges arising from heterogeneity, application and infrastructure dynamism, complex integration needs, and software fragmentation [12]. While quantum-HPC systems share these difficulties, they substantially amplify them due to additional complexities unique to quantum computing. For example, leveraging quantum-HPC systems fully requires integrating multiple new types of infrastructure and support for new parallelism and integration patterns, e.g., circuit-level parallelism across multiple QPUs, tightly-coupled GPU-QPU integration for error correction, and loosely-coupled integration of variational quantum algorithms (VQA). Additional new task constraints, including qubit counts, fidelity thresholds, and classical-quantum resource co-location, must also be considered for scheduling and execution. Further, the rapidly evolving and often fragmented quantum hardware, software libraries [13, 14, 15], and middleware [16, 17, 18] landscape makes integration and scaling difficult.
Quantum-HPC Middleware: Modern workflow systems are increasingly shifting from monolithic designs to modular architectures that separate concerns across distinct layers [12]. Following this principle, we introduce Pilot-Quantum, a modular middleware for resource, workload, and task management that provides an application-level framework for orchestrating hybrid quantum-classical workloads across CPUs, GPUs, and QPUs. Pilot-Quantum addresses the key challenges of hybrid execution through multi-level scheduling, enabling adaptive resource allocation tailored to application needs and efficient use of costly and heterogeneous resources. Pilot-Quantum builds upon the Pilot-Abstraction [19, 20], a conceptual model for quantum-HPC middleware [21] and an analysis of execution motifs [22]. It supports integration with high-level frameworks like Qiskit [14, 23] and PennyLane [15] and low-level runtimes like cuQuantum [24]. This design enables scalable, efficient execution of complex hybrid workflows in rapidly evolving quantum-HPC environments.
@@
This paper is structured as follows: We explore quantum software and middleware state-of-the-art in section II. The primary contribution is Pilot-Quantum, an open-source middleware [25] addressing resource, workload, and task management challenges, introduced in section III. Section IV demonstrates its capabilities by extending quantum mini apps [22] introducing, e.g., a circuit cutting and quantum machine learning (QML) mini-apps. We conclude with a qualitative comparison of quantum middleware systems with Pilot-Quantum and future work in section V.
@@ II Background and Related Work
This section analyzes quantum-classical software and middleware using a conceptual architecture model [21], quantum-HPC integration modes and execution motifs [22].
@@ II-A Conceptual Middleware Architecture
In [21], we define a conceptual architecture for a high-level system design of quantum-HPC middleware systems that explains the main components and their interactions. It consists of four distinct layers, denoted as $L1$, $L2$, $L3$, $L4$. $L4$ is the workflow layer, which encapsulates application semantics and logical dependencies between the workflow stages and offers a high-level abstraction for developing the quantum-HPC workflows. The workload layer $L3$ manages the workload, a set of interdependent, quantum, and classical tasks emitted by $L4$. It optimizes the workload execution by partitioning tasks and assigning them to the available resources utilizing parallelism (e.g., ensemble and task parallelism).
@@ II-B Quantum-HPC Integration Modes and Motifs
Understanding common quantum-HPC integration modes and motifs (i.e., recurring execution patterns) between quantum-HPC components [22] is critical to designing effective middleware systems.
@@
are categorized as Quantum-x-HPC or HPC-x-Quantum based on the strength of coupling between quantum and classical tasks [7, 21]. Three modes are identified: HPC-for-Quantum, with strong coupling optimizing quantum-classical interactions (e. g., error mitigation [26], error correction [27, 28], dynamic circuits [29]); Quantum-in-HPC emphasizes loosely-coupled parallelism between classical and quantum resources found in VQA [30] to accelerate HPC applications; and Quantum-about-HPC, with loose coupling of quantum elements into HPC applications for pre- and post-processing [31, 32].
@@ II-C Quantum Software Layers
frameworks typically provide quantum circuit model-based abstraction for describing quantum programs. Examples are Qiskit [14, 23], PennyLane [15], Cirq [33], low-level frameworks like CUDA-Q and AWS Braket [34], and high-level programming libraries, such as Qrisp [35], Silq [36], and Q# [37]. At the $L4+$ levels, domain-specific programming libraries include Qiskit Optimization [38], Qiskit Nature [39], and OpenFermion [40].
@@
like Qiskit Runtime [41] and Braket Hybrid Jobs [42, 43] provide session and decorator abstractions, respectively, to co-allocate heterogeneous resources for hybrid applications. However, new execution patterns, such as the combination of quantum and machine learning capabilities, require new abstractions, including auto differentiation of quantum circuits (e. g., using PennyLane [15]) or training classical models using quantum task outputs [44]. Some frameworks like XACC [45], QCOR [46], and CUDA-Q [18] use a kernel-based programming model with compile toolchains to manage quantum-classical function interactions. For classical simulation of quantum systems, acceleration frameworks, e. g., cuQuantum [24] are critical for performance and scale.
@@
At the $L1$ layer, many quantum frameworks provide device abstractions and plugins for accessing quantum devices, such as Qiskit backends [47], Qiskit Runtime [41], and PennyLane plugins [48]. Intermediate representations like OpenQASM [49] and QIR [50, 51] aim to bridge the gap between hardware platforms and high-level programming constructs. Transpilers and compilers are used to transform a circuit into a hardware-optimized form. They are provided by full-stack frameworks like Qiskit and specialized frameworks like Tket [52], BQSKit [53], and Staq [54] transpilers.
@@
Resource management and scheduling in quantum-HPC environments occur across multiple levels and time scales [8, 55], involving integration of coarse-grained job-level allocation (e. g., GPUs) with QPU allocations and fine-grained, application-level scheduling [20, 56]. Existing HPC resource management frameworks often assume exclusive resource access, leading to resource fragmentation and underutilization [57]. As algorithms scale across heterogeneous resources (e. g., CPUs, GPUs, QPUs), such resource management will result in suboptimal performance and utilization. Pilot-Quantum specifically addresses the need for an application-level resource management by balancing system and application objectives.
@@
at the task and workload layer ($L3$ and $L4$) are needed to orchestrate quantum workloads across integrated quantum-HPC systems. QML workloads, for instance, involve heterogeneous classical-quantum tasks for gradient computation, backpropagation, optimization, and circuit execution, requiring a systematic approach to workload and task management.
@@ II-D Middleware for Resource, Workload and Task Management
Many quantum middleware systems emerged recently including, QFaaS [58], Tierkreis [59], Qibo [60], Quantum Framework (QFw) [61], Quantum Brilliance [62], Rasqal [63] and Zapata’s Orquestra [64]. We particularly investigate Qiskit Serverless [16], Covalent [17], and CUDA-Q [18] systems.
II-D1 Qiskit Serverless [65, 66, 16, 67]
is a middleware for managing resources and workloads in hybrid quantum-classical infrastructures while supporting long-running applications and scaling across distributed resources (CPU, GPU, QPU). It extends cloud-based quantum hardware access to classical computing environments at the $L2$ and $L3$ layers, utilizing Apache Ray [68] to execute hybrid workloads.
@@ II-D1 Qiskit Serverless [65, 66, 16, 67]
II-D2 Covalent [17, 71]
orchestrates distributed computing workflows across Cloud-HPC resources. Its SDK provides $L4$ level abstraction using electron and lattice decorators for defining tasks and workflows (i. e., DAG of electrons), where quantum tasks are defined using qelectron decorator. The Dispatch service at the $L3$ layer accepts workloads (i. e., ready-to-run tasks), maps them to suitable resources, and dispatches them to appropriate backends.
@@ II-D3 CUDA-Q
[18] supports tight integrations within heterogeneous computing environments. It employs an LLVM compiler toolchain [74] to convert frontend code in C++ or Python to different heterogeneous backends, such as CPUs, CUDA-enabled GPUs, and QPUs. It utilizes distinct intermediate representations for each processor type (e. g., QIR for QPUs and NVVM IR for GPUs), allowing for targeted optimizations. The CUDA-Q runtime, utilizes MPI for parallel processing and CUDA as a low-level runtime. While particularly suitable for HPC-for-Quantum tight integration, CUDA-Q also supports Quantum-for-HPC modes, e. g., for VQAs. At $L4$, it provides a circuit-based programming model and a kernel API for offloading circuits to hardware or simulated QPUs. Application-level task management uses an MPI and CUDA-Runtime on the $L2$/$L1$ levels.
@@ III Pilot-Quantum Design and Architecture
Pilot-Quantum manages resources, workloads, and tasks across heterogeneous classical and quantum infrastructures for quantum applications and workflows. It is an implementation of the Pilot Abstraction [19, 20], a conceptual abstraction for application-level scheduling [19, 75]. It uses placeholder jobs for application-level resource control, enabling flexible and adaptive scaling of resources and workloads.
Abstraction for Resource Management: Pilot-Quantum enables applications to allocate resources (i.e., nodes, cores, GPUs, and QPUs) and execute task-based workloads on them. It provides application-level control over resources, enabling adaptive optimization of resource allocations to workloads and vice versa. This allows the application to dynamically acquire or release resources at runtime, such as optimizing the mix of classical and quantum resources for a specific workflow step. It supports typical workloads in quantum workflows, such as running circuit ensembles, decomposing larger circuits via circuit cutting, and handling data-parallel QML workloads.
Easy Deployment: Pilot-Quantum simplifies workflow deployment across distributed, heterogeneous infrastructures, including non-local quantum clouds, HPC MPI/CUDA environments for distributed state vector simulation, and loosely coupled workflow systems. It can also integrate with resource-specific optimization tools like QPU transpilers.
@@ III-A Requirements and Design Objective
Infrastructure Heterogeneity: The system must address software and hardware heterogeneity, encompassing a range of QPUs like superconducting qubits, ion traps, neutral atoms, and simulated QPUs using MPI or CUDA. This requires compatibility with diverse frameworks and tools, such as transpilers, compilers, and runtimes, including simulators (e. g., state vector and tensor network simulations with MPI or CUDA) and machine learning runtimes like PyTorch and JAX.
@@
Resource management in quantum-classical computing is challenging due to diverse resource types and application demands. Dynamic allocation during workflow execution is crucial to address application and infrastructure dynamism when handling tasks and dependencies across heterogeneous platforms. Multi-level scheduling enables aligning application and system objectives, efficient resource allocation, workload partitioning, and task management strategies to optimize application performance and system utilization.
Abstractions: Integrating quantum computing into HPC workflows requires flexible abstractions that hide quantum complexity while supporting various parallelism forms (e.g., task parallelism for circuits, distributed state vector simulations using MPI and cuQuantum, and differentiable computing for QML). These should facilitate efficient pipeline creation for end-to-end quantum applications with pre- and post-processing integrations. The abstractions must offer seamless interoperability with quantum and HPC infrastructures, ensuring compatibility with frameworks like CUDA for GPU-accelerated tasks and MPI for distributed computing.